# uncompress

### v2

@@ 2.3. Singularities of pairs
In this paper, we will use standard notions and results from minimal model program; cf. [KM98, BCHM10]. Here we collect some of the most fundamental definitions for clarification. A pair $(X,B)$ consists of a normal quasi-projective variety $X$ and an $\mathbb{R}$-divisor $B$ with coefficients in $[0,1]$ such that $K_{X}+B$ is $\mathbb{R}$-Cartier; in this case, we say that $B$ is a boundary. Let $W\to X$ be a log resolution of a pair $(X,B)$, and let $K_{W}+B_{W}$ be the pullback of $K_{X}+B$. Denote by $\mu_{D}B_{W}$ the coefficient of $B_{W}$ at a prime divisor $D$ on $W$, then the log discrepancy of $D$ with respect to $(X,B)$ is $1-\mu_{D}B_{W}$, and it is denoted by $a(D,X,B)$. We say that $(X,B)$ is lc (respectively, klt, respectively, $\epsilon$-lc) if $a(D,X,B)$ is $\geq 0$ (respectively, >0>0, respectively, $\geq\epsilon$) for every divisor $D$ on an arbitrary log resolution $W\to X$. Note that $a(D,X,B)$ can also be defined in the same way when the coefficients of $B$ do not belong to $[0,1]$ as long as $K_{X}+B$ is $\mathbb{R}$-Cartier.
Let $(X,B)$ be an lc pair. An lc place of $(X,B)$ is a prime divisor on some birational model of $X$ such that $a(D,X,B)=0$. An lc centre is the image on $X$ of an lc place.
For dlt pairs, we refer the readers to [KM98, §2.3] for the basic definitions and to [KK10] for the definition of dlt modifications (in particular, see [KK10, Theorem 3.1] for the existence of $\mathbb{Q}$-factorial dlt modifications).
For the basic definitions and properties of generalised pairs, we recommend the readers to [BZ16, §4] and [Bir19, §2]. For the definition and existence of $\mathbb{Q}$-factorial generalised dlt modifications (also called $\mathbb{Q}$-factorial generalised dlt models); see [Bir19, §2.13].
Singularities of pairs, such as lc singularities, can also be defined for demi-normal schemes, and the corresponding pairs with lc property are called slc pairs. We refer to [Kol13, Chapter 5] for more details about slc pairs.
The following result can be shown easily by taking common log resolutions and pulling back the pairs, and the proof is left to readers.
@@
Let $X$ and $Y$ be normal quasi-projective varieties over a variety $Z$, and let $\phi\colon Y\dashrightarrow X$ be a birational map. Let $B$ be a $\mathbb{Q}$-divisor on $X$ such that $K_{X}+B$ is $\mathbb{Q}$-Cartier and $F$ a divisor over $X$. Define a divisor $B_{Y}$ on $Y$ by $K_{Y}+B_{Y}=\phi^{*}(K_{X}+B)$. Assume that $K_{X}+B\sim_{\mathbb{Q}}0/Z$. Then the log discrepancy $a(F,X,B)$ is equal to $a(F,Y,B_{Y})$.
@@ 2.5. Fano fibrations
Let $(X,B)$ be an lc pair over a variety $Z$. We say that $(X,B)$ is log Fano over $Z$ if $(K_{X}+B)$ is ample over $Z$. If $Z$ is a point, then $(X,B)$ is called a log Fano pair. In this case, if $B=0$, then $X$ is called a Fano variety. A morphism $(X,B)\to Z$ from an lc pair $(X,B)$ is a log Fano fibration if $(X,B)$ is log Fano over $Z$ and the underlying morphism $X\to Z$ is a contraction. A klt log Fano fibration $(X,B)\to Z$ is a log Fano fibration with $(X,B)$ klt. In this case, if $B=0$, we call $X\to Z$ a klt Fano fibration; see [BL21, §3].
@@ 2.6. Complements
We define complements as in [Sho00]; see also [Bir19, §2.18]. Let $(X,B)$ be a pair where $B$ is a boundary, and let $X\to Z$ be a contraction. Let $T=\left\lfloor{B}\right\rfloor$ and $\Delta=B-T$, and $n\in\mathbb{N}$ a natural number. An $n$-complement of $K_{X}+B$ over a point $z\in Z$ is of the form $K_{X}+B^{+}$ such that over some neighbourhood of $z$ we have the following properties:
@@ 2.7. Couples
A couple $(X,D)$ consists of a quasi-projective variety $X$ and a reduced Weil divisor $D$ on $X$. This is more general than the definition given in [Bir19, §2.19] because we are not assuming $X$ to be normal nor projective. Also note that a couple is not necessarily a pair in the sense that we are not assuming $K_{X}+D$ to be $\mathbb{R}$-Cartier. In this paper, we often consider a couple $(X,D)$ equipped with a surjective projective morphism to a variety $Z$ in which case we often denote the couple as $(X/Z,D)$ or $(X,D)\to Z$.
@@ 2.8. Strata and log smooth morphisms
We follow the conventions as in [HMX18, §2.1]. Let $(X,D)$ be a couple. The strata of $(X,D)$ are the irreducible components of the intersections
@@
of components of $D$, where $I=\{i_{1},\dots,i_{r}\}$ is a subset of the indices, including the empty intersection $X=D_{\emptyset}$. Every irreducible component in the strata of $(X,D)$ is called a stratum of $(X,D)$. If $(X,B)$ is a pair, then the strata of $(X,B)$ are the strata of the underlying couple $(X,D)$ where $\operatorname{Supp}D=\operatorname{Supp}B$.
If we are given a couple $(X,D)$ over a variety $T$, then we say that $(X,D)$ is log smooth over $T$ if $(X,D)$ has simple normal crossings and the strata of $(X,D)$ are smooth over $T$. If $(X,B)$ is a pair over a variety $U$, we say that $(X,B)$ is log smooth over $U$ if the underlying couple $(X,D)$ is log smooth over $U$ where $\operatorname{Supp}D=\operatorname{Supp}B$.
@@ 2.9. Bounded and birationally bounded families of couples
We say that a set $\mathcal{Q}$ of couples $(X,D)$ is bounded if there exist finitely many projective morphisms $V_{i}\to T_{i}$ of varieties and reduced divisors $C_{i}$ on $V_{i}$ such that for each $(X,D)\in\mathcal{Q}$, there exist an $i$, a closed point $t\in T_{i}$, and an isomorphism of couples $\phi\colon(V_{i,t},C_{i,t})\to(X,D)$, where $V_{i,t}$ and $C_{i,t}$ are the fibres over $t$ of the morphisms $V_{i}\to T_{i}$ and $C_{i}\to T_{i}$ respectively. In particular, if $D=0$ for all $(X,D)\in\mathcal{Q}$, we say the family $\mathcal{Q}$ consisting of projective varieties $X$ is bounded.
Let $\mathcal{P}$ be a set of projective varieties. We say that $\mathcal{P}$ is birationally bounded if there is a bounded family $\mathcal{Q}$ of projective varieties such that for every $X\in\mathcal{P}$ there exists a birational map $X\dashrightarrow Y$ to a member $Y$ in $\mathcal{Q}$. Note that our definition of birational boundedness is weaker than [Bir19, §2.19].
@@
We say a set $\mathcal{S}$ of projective pairs $(X,B)$ is bounded if there exist finitely many projective morphisms $Z_{i}\to U_{i}$, where $U_{i}$ is smooth, $Z_{i}$ is flat over $U_{i}$, and pairs $(Z_{i},\Sigma_{i})$, where the support of $\Sigma_{i}$ contains neither a component of a fibre nor a codimension one singular point of any fibre, such that for every $(X,B)\in\mathcal{S}$, there exist an $i$, a closed point $u\in U_{i}$, and an isomorphism of pairs $\psi\colon(Z_{i,u},\Sigma_{i,u})\to(X,B)$, where $Z_{i,u}$ and $\Sigma_{i,u}$ are the fibres over $u$ of the morphisms $Z_{i}\to U_{i}$ and $\Sigma_{i}\to U_{i}$ respectively. In particular, the coefficients of $B$ belong to a finite set; see [HMX18, page 873].
The following lemma is well-known and we omit its proof.
@@
Let $d$ be a natural number. Let $(X,B+M)$ be a projective generalised lc generalised pair with $\dim X=d$ such that $L=K_{X}+B+M$ is a nef and big integral Cartier divisor. Then $\left|mL\right|$ defines a birational map for some $m\in\mathbb{N}$ depending only on $d$.
@@
Taking a $\mathbb{Q}$-factorial dlt modification $f\colon Y\to X$ of $(X,B+M)$ (see [BZ16, Lemma 4.5]), we have that
$K_{Y}+B_{Y}+M_{Y}=f^{*}(K_{X}+B+M)$
where $B_{Y}$ is the sum of $f^{-1}_{*}B$ and the reduced $f$-exceptional divisors and $(Y,B_{Y}+M_{Y})$ is a $\mathbb{Q}$-factorial generalised dlt generalised pair. In particular, $(Y,0)$ is $\mathbb{Q}$-factorial and klt. Fix a small rational number $0<\epsilon\ll 1$. Let $g\colon W\to Y$ be a birational morphism extracting the divisors $F$ whose log discrepancies with respect to $(Y,0)$ are $\leq\epsilon$, then $W$ is $\epsilon$-lc. Write
@@
Then the coefficients of $E$ are $\geq 1-\epsilon$. Now, we can write
$(f\circ g)^{*}L-K_{W}=E+g^{*}(B_{Y}+M_{Y})$
which is clearly pseudo-effective. Since $(f\circ g)^{*}L$ is also a nef and big integral Cartier divisor, we can apply [Bir23, Theorem 1.1] to conclude. ∎
@@ 2.12. Relative degree
Let $f\colon X\to Z$ be a surjective projective morphism of quasi-projective varieties and let $A$ be a Cartier divisor on $X$. For a Weil divisor $D$ on $X$ we define the relative degree of $D$ with respect to $A$ as $\deg_{A/Z}D\coloneqq(D|_{F})\cdot(A|_{F})^{n-1}$ where $F$ is a general fibre of $f$ and $n=\dim F$. It is clear that this is a generic property, so the vertical/Z/Z components of $D$ do not contribute to the relative degree. Note that $F$ may not be irreducible.

### v3

@@ 2.3. Singularities of pairs
In this paper, we will use standard notions and results from minimal model program; see [12, 36]. Here we collect some of the most fundamental definitions for clarification. A pair $(X,B)$ consists of a normal quasi-projective variety $X$ and an $\mathbb{R}$-divisor $B$ with coefficients in $[0,1]$ such that $K_{X}+B$ is $\mathbb{R}$-Cartier; in this case, we say that $B$ is a boundary. Let $W\to X$ be a log resolution of a pair $(X,B)$, and let $K_{W}+B_{W}$ be the pullback of $K_{X}+B$. Denote by $\mu_{D}B_{W}$ the coefficient of $B_{W}$ at a prime divisor $D$ on $W$. Then the log discrepancy of $D$ with respect to $(X,B)$ is $1-\mu_{D}B_{W}$, and it is denoted by $a(D,X,B)$. We say that $(X,B)$ is lc (respectively, klt, respectively, $\epsilon$-lc) if $a(D,X,B)$ is $\geq 0$ (respectively, >0>0, respectively, $\geq\epsilon$) for every prime divisor $D$ on an arbitrary log resolution $W\to X$. Note that $a(D,X,B)$ can also be defined in the same way when the coefficients of $B$ do not belong to $[0,1]$ as long as $K_{X}+B$ is $\mathbb{R}$-Cartier.
Let $(X,B)$ be an lc pair. An lc place of $(X,B)$ is a prime divisor $D$ on some birational model of $X$ such that $a(D,X,B)=0$. An lc centre is the closure of the image of an lc place on $X$.
For dlt pairs, we refer the readers to [36, §2.3] for the basic definitions. Let $(X,B)$ be an lc pair. A $\mathbb{Q}$-factorial dlt model of $(X,B)$ is a $\mathbb{Q}$-factorial dlt pair $(Y,B_{Y})$ admitting a projective birational morphism $\pi\colon Y\to X$ such that $K_{Y}+B_{Y}=\pi^{*}(K_{X}+B)$ and that every $\pi$-exceptional prime divisor is an irreducible component of $\left\lfloor{B_{Y}}\right\rfloor$; cf. [13, §5.1]. In particular, see [35, Theorem 3.1] for the existence of $\mathbb{Q}$-factorial dlt models.
For the basic definitions and properties of generalised pairs, we refer the readers to [7, §2] and [14, §4]. For the definition and existence of $\mathbb{Q}$-factorial generalised dlt models, see [7, §2.13].
Singularities of pairs, such as lc singularities, can also be defined for demi-normal schemes, and the corresponding pairs with lc property are called slc pairs. We refer the readers to [34, Chapter 5] for more details about slc pairs.
@@
Let $X$ and $Y$ be normal varieties over a variety $Z$, and let $\phi\colon Y\dashrightarrow X$ be a birational map over $Z$. Let $B$ be an $\mathbb{R}$-divisor on $X$ such that $K_{X}+B$ is $\mathbb{R}$-Cartier, and let $B_{Y}$ be the $\mathbb{R}$-divisor on $Y$ defined by $K_{Y}+B_{Y}=\phi^{*}(K_{X}+B)$. Let $F$ be a prime divisor over $X$. Assume that $K_{X}+B\sim_{\mathbb{R}}0/Z$. Then the log discrepancy $a(F,X,B)$ is equal to $a(F,Y,B_{Y})$.
Proof.
Denote by $f,g$ the morphisms $X\to Z,Y\to Z$ respectively. Let $p\colon W\to X,q\colon W\to Y$ be a common resolution of $X,Y$ such that $f\circ p=g\circ q$. Then $\phi^{*}(K_{X}+B)$ is defined to be the $\mathbb{R}$-divisor $q_{*}p^{*}(K_{X}+B)$. It is easy to see that $\phi^{*}(K_{X}+B)$ does not depend on the choice of $p,q$, hence the $\mathbb{R}$-divisor $B_{Y}\coloneqq\phi^{*}(K_{X}+B)-K_{Y}$ is well-defined. Replacing $W$ by higher resolutions, we can assume that $p$ is a log resolution of $(X,B)$, $q$ is a log resolution of $(Y,B_{Y})$, and $F$ is a prime divisor on $W$.
By assumption, there is an $\mathbb{R}$-Cartier $\mathbb{R}$-divisor $N$ on $Z$ such that $K_{X}+B=f^{*}N+\sum a_{i}\operatorname{Div}_{X}(\alpha_{i})$, where $a_{i}\in\mathbb{R}$, $\alpha_{i}\in K(X)$, and $\operatorname{Div}_{X}(\alpha_{i})$ is the principal divisor on $X$ associated to $\alpha_{i}$. As $\phi$ is birational, we can view each $\alpha_{i}$ as a rational function on $Y$ (also on $W$). Then it is evident that
$p^{*}(K_{X}+B)=(f\circ p)^{*}N+\sum a_{i}\operatorname{Div}_{W}(\alpha_{i})=(g\circ q)^{*}N+\sum a_{i}\operatorname{Div}_{W}(\alpha_{i}),$
which implies $K_{Y}+B_{Y}=g^{*}N+\sum a_{i}\operatorname{Div}_{Y}(\alpha_{i})$. In particular, $K_{Y}+B_{Y}$ is $\mathbb{R}$-Cartier, and $q^{*}(K_{Y}+B_{Y})=p^{*}(K_{X}+B)$. Then we can conclude that $a(F,X,B)=a(F,Y,B_{Y})$. ∎
@@ 2.5. Fano fibrations
Let $(X,B)$ be an lc pair over a variety $Z$. We say that $(X,B)$ is log Fano over $Z$ if $(K_{X}+B)$ is ample over $Z$. If $Z$ is a point, then $(X,B)$ is called a log Fano pair. In this case, if $B=0$, then $X$ is called a Fano variety.
A morphism $(X,B)\to Z$ from an lc pair $(X,B)$ is a log Fano fibration if $(X,B)$ is log Fano over $Z$ and the underlying morphism $X\to Z$ is a contraction; moreover, if $B=0$, we call $X\to Z$ a Fano fibration. A klt log Fano fibration $(X,B)\to Z$ is a log Fano fibration with $(X,B)$ klt. In this case, if $B=0$, we call $X\to Z$ a klt Fano fibration; see [13, §3].
@@ 2.6. Complements
We define complements as in [41]; see also [7, §2.18]. Let $(X,B)$ be a pair, and let $X\to Z$ be a contraction. Let $T=\left\lfloor{B}\right\rfloor$ and $\Delta=B-T$, and $n\in\mathbb{N}$ a natural number. An $n$-complement of $K_{X}+B$ over a point $z\in Z$ is of the form $K_{X}+B^{+}$ such that over some neighbourhood of $z$, we have the following properties:
@@ 2.7. Couples
A couple $(X,D)$ consists of a quasi-projective variety $X$ and a reduced Weil divisor $D$ on $X$. This is more general than the definition given in [7, §2.19] because we are not assuming $X$ to be normal or projective. Also note that a couple is not necessarily a pair in the sense that we are not assuming $K_{X}+D$ to be $\mathbb{R}$-Cartier. In this paper, we often consider a couple $(X,D)$ equipped with a surjective projective morphism to a variety $Z$ in which case we often denote the couple as $(X/Z,D)$ or $(X,D)\to Z$; if $Z$ is a point, we call $(X,D)$ a projective couple.
@@ 2.8. Strata and log smooth morphisms
We follow the conventions as in [26, §2.1]. Let $(X,D)$ be a couple. The strata of $(X,D)$ are the irreducible components of the intersections
@@
of irreducible components of $D$, where $I=\{i_{1},\dots,i_{r}\}$ is a subset of the indices, including the empty intersection $X=D_{\emptyset}$. Every irreducible component in the strata of $(X,D)$ is called a stratum of $(X,D)$. If $(X,B)$ is a pair, then the strata of $(X,B)$ are the strata of the underlying couple $(X,D)$, where $\operatorname{Supp}D=\operatorname{Supp}B$.
If we are given a couple $(X,D)$ over a variety $T$, then we say that $(X,D)$ is log smooth over $T$ if $(X,D)$ has simple normal crossings and the strata of $(X,D)$ are smooth over $T$. If $(X,B)$ is a pair over a variety $U$, we say that $(X,B)$ is log smooth over $U$ if the underlying couple $(X,D)$ is log smooth over $U$, where $\operatorname{Supp}D=\operatorname{Supp}B$.
@@ 2.9. Bounded and birationally bounded families of couples
We say that a set $\mathcal{Q}$ of projective couples $(X,D)$ is bounded if there exist finitely many projective morphisms $V_{i}\to T_{i}$ of varieties and reduced divisors $C_{i}$ on $V_{i}$ such that for each $(X,D)\in\mathcal{Q}$, there exist an $i$, a closed point $t\in T_{i}$, and an isomorphism of couples $\phi\colon(V_{i,t},C_{i,t})\to(X,D)$, where $V_{i,t}$ and $C_{i,t}$ are the fibres over $t$ of the morphisms $V_{i}\to T_{i}$ and $C_{i}\to T_{i}$ respectively. In particular, if $D=0$ for all $(X,D)\in\mathcal{Q}$, we say the family $\mathcal{Q}$ consisting of projective varieties $X$ is bounded.
Let $\mathcal{P}$ be a set of projective varieties. We say that $\mathcal{P}$ is birationally bounded if there is a bounded family $\mathcal{Q}$ of projective varieties such that for every $X\in\mathcal{P}$, there exists a birational map $X\dashrightarrow Y$ to a member $Y$ in $\mathcal{Q}$. Note that our definition of birational boundedness is weaker than [7, §2.19].
@@
We say a set $\mathcal{S}$ of projective pairs $(X,B)$ is bounded if there exist finitely many projective morphisms $Z_{i}\to U_{i}$, where $U_{i}$ is smooth, $Z_{i}$ is flat over $U_{i}$, and pairs $(Z_{i},\Sigma_{i})$, where the support of $\Sigma_{i}$ contains neither an irreducible component of a fibre nor a codimension one singular point of any fibre, such that for every $(X,B)\in\mathcal{S}$, there exist an $i$, a closed point $u\in U_{i}$, and an isomorphism of pairs $\psi\colon(Z_{i,u},\Sigma_{i,u})\to(X,B)$, where $Z_{i,u}$ and $\Sigma_{i,u}$ are the fibres over $u$ of the morphisms $Z_{i}\to U_{i}$ and $\Sigma_{i}\to U_{i}$ respectively. In particular, the coefficients of $B$ belong to a finite set; cf. [26, page 873].
@@
Proof.
Let $(X,B)\in\mathcal{Q}$. By assumption, there exist finitely many projective morphisms $Z_{i}\to U_{i}$ of varieties, where $U_{i}$ is smooth, $Z_{i}$ is flat over $U_{i}$, and pairs $(Z_{i},\Sigma_{i})$ such that there exist an $i$, a closed point $u\in U_{i}$, and an isomorphism of pairs $(Z_{i,u},\Sigma_{i,u})\to(X,B)$. Let $\pi_{i}\colon W_{i}\to Z_{i}$ be a log resolution of $(Z_{i},\Sigma_{i})$, and write
$K_{W_{i}}+\Gamma_{i}=\pi_{i}^{*}(K_{Z_{i}}+\Sigma_{i}).$
Then there exists a nonempty open subset $V_{i}\subseteq U_{i}$ such that for every closed point $v\in V_{i}$, $(Z_{i,v},\Sigma_{i,v})$ is a pair, the induced morphism of fibres $\pi_{i,v}\colon W_{i,v}\to Z_{i,v}$ is a log resolution of $(Z_{i,v},\Sigma_{i,v})$, and that $\pi_{i,v}^{*}(K_{Z_{i,v}}+\Sigma_{i,v})$ is the restriction of $\pi_{i}^{*}(K_{Z_{i}}+\Sigma_{i})$ to $W_{i,v}$. The complement $U_{i}\setminus V_{i}$ is a union of varieties with dimensions strictly less than $\dim U_{i}$. By induction on $\dim U_{i}$, we can assume that $u\in V_{i}$. Then every lc centre of $(Z_{i,u},\Sigma_{i,u})$ is the image of a stratum of $\left\lfloor{\Gamma_{i,u}}\right\rfloor$ via $\pi_{i,u}\colon W_{i,u}\to Z_{i,u}$. Thus, there exist finitely many integral subschemes of $Z_{i}\times_{U_{i}}V_{i}$ such that every lc centre of $(X,B)$ is isomorphic to the fibre of one of these subschemes over $u\in V_{i}$. Hence the set of lc centres of $(X,B)$ is bounded. ∎
@@
Let $d$ be a natural number. Let $(X,B+M)$ be a projective generalised lc generalised pair with $\dim X=d$ such that $L=K_{X}+B+M$ is a nef and big Cartier divisor. Then $\left|mL\right|$ defines a birational map for some $m\in\mathbb{N}$ depending only on $d$.
@@
Taking a $\mathbb{Q}$-factorial generalised dlt model $f\colon Y\to X$ of $(X,B+M)$ (see [14, Lemma 4.5]), we have that
$K_{Y}+B_{Y}+M_{Y}=f^{*}(K_{X}+B+M),$
where $B_{Y}$ is the sum of $f^{-1}_{*}B$ and the reduced $f$-exceptional divisors, and $(Y,B_{Y}+M_{Y})$ is a $\mathbb{Q}$-factorial generalised dlt generalised pair. In particular, $(Y,0)$ is $\mathbb{Q}$-factorial and klt. Fix a small rational number $0<\epsilon\ll 1$. Let $g\colon W\to Y$ be a birational morphism extracting the divisors whose log discrepancies with respect to $(Y,0)$ are $\leq\epsilon$. Then $W$ is $\epsilon$-lc. Write
@@
Then the coefficients of $E$ are $\geq 1-\epsilon$. Now we can write
$(f\circ g)^{*}L-K_{W}=E+g^{*}(B_{Y}+M_{Y}),$
which is clearly pseudo-effective. Since $(f\circ g)^{*}L$ is also a nef and big Cartier divisor, we can apply [9, Theorem 1.1] to conclude. ∎
@@ 2.12. Relative degree
Let $f\colon X\to Z$ be a surjective projective morphism of quasi-projective varieties, and let $A$ be a $\mathbb{Q}$-Cartier divisor on $X$. For a Weil divisor $D$ on $X$, we define the relative degree of $D$ over $Z$ with respect to $A$ as $\deg_{A/Z}D\coloneqq(D|_{F})\cdot(A|_{F})^{n-1}$, where $F$ is a general fibre of $f$ and $n=\dim F$. It is clear that this is a generic property, so the vertical/Z/Z irreducible components of $D$ do not contribute to the relative degree. Note that $F$ may not be irreducible: by a general fibre, we mean the fibre over a general closed point of $Z$. In practice, we take $A$ to be ample over $Z$; see [11, §2.2].